# Understanding and Patching Compositional Reasoning in LLMs

### v1

@@ Appendix A Related Works on Mechanistic Interpretability and Model Editing
Mechanistic Interpretability, interpreting inner workings of LLMs, is drawing an increasing attention of NLP researchers. Logit Lens nostalgebraist (2020) is proposed to interpret hidden states at the middle layers of LLMs via projecting them to the output vocabulary space with the LM head. Subsequent works Geva et al. (2021, 2022); Dar et al. (2023); Katz and Belinkov (2023) further explain how LLMs build precise next token predictions. Another line of MI works focus on inspecting factual knowledge encoded in the LLMs: they first locate such factual knowledge in pretrained LLMs Dai et al. (2022); Geva et al. (2023); Li et al. (2023b) and then edit them through updating a small set of parameters of LLMs Meng et al. (2022, 2023); Hase et al. (2023), which is so-called “locate-then-edit” model editing Ju and Zhang (2023). In this paper, we shed light on the mechanism of compositional reasoning on factual knowledge and borrow the idea from “locate-then-edit” model editing to correct compositional reasoning failures of LLMs.
@@ D.3 Important Module Locating
We implement our locating method on the basis of Causal Tracing Meng et al. (2022). Following  Meng et al. (2022)’s implementation, we also use a “window” intervention(a few layers before and after the intervened layer). In their original codebase, they set window size to be 10. In our experiments: we find that setting window size to be 2 is enough for us to effectively locate important modules.
@@ D.4 Model Editing:CREME
The edit objective is the MHSA output matrix of $l$-th layer $W_{O}^{l}$. $W_{O}^{l}\in\mathbb{R}^{d\times d}$ operates as a key-value store for a set of vector keys $K=[k_{1}|k_{2}|...]$ and corresponding vector values $V=[v_{1}|v_{2}|...]$, by solving $(W_{O}^{l})^{T}K=V$. For a given compositional two-hop query and its corresponding second-hop query, we run the LLM twice: once with the compositional query and once with the second-hop query. We denote that: in the first pass with compositional query, the input of $W_{O}^{l}$ at the last token position is $k_{*}\in\mathbb{R}^{d\times 1}$; in the second pass with the corresponding second-hop query, the output of $W_{O}^{l}$ at the last token position is $v_{*}\in\mathbb{R}^{d\times 1}$. In practice, when calculating $k_{*}$ and $v_{*}$, we prepend tens of random tokens to the compositional query and the corresponding second-hop query to mimic context environments, and get multiple input vectors and output vectors. Then we average input vectors and output vectors of different context environment to get $k_{*}$ and $v_{*}$, respectively. The edited matrix is: $\hat{W_{O}^{l}}=W_{O}^{l}+(C^{-1}k_{*})^{T}\Lambda^{T}$ where $C=KK^{T}$ is a constant to estimate the uncentered covariance of $k$ (with a sample Wikipedia of text) and $\Lambda=(v_{*}-(W_{O}^{l})^{T}k_{*})/(C^{-1}k_{*})^{T}k_{*}$.
@@ Appendix E Additional Results
Testing type Input Prediction w.o. CREME Prediction w. CREME
Hasty Answer II
Incomplete Reasoning
Table 5: Case study for correcting the (1) Hasty Answer II error: the original input (used for correcting) is “The country that the creator of C. Auguste Dupin belongs to is”. The original prediction is “France” (Reference: C. Auguste Dupin is French, while his creator Edgar Allan Poe. is American.); and (2) Incomplete Reasoning error: the original input (used for correcting) is “The capital of the country that Sven Väth is a citizen of is”. The original prediction is “Germany” (Reference: Berlin.).
@@ E.2 Intervention Results
(a) LLaMA-2-7B
(b) OpenAlpaca-3B
Figure 11: Intervention experiment: Brighter color indicates the intervention effect is more significant. In each subfigure, the upper row refers to the experiment group and the lower row refers to the comparison group.

### v2

@@ Appendix A Related Works on Mechanistic Interpretability and Model Editing
Mechanistic Interpretability, interpreting inner workings of LLMs, is drawing an increasing attention of NLP researchers. Logit Lens nostalgebraist (2020) is proposed to interpret hidden states at the middle layers of LLMs via projecting them to the output vocabulary space with the LM head. Subsequent works Geva et al. (2021, 2022); Dar et al. (2023); Katz and Belinkov (2023) further explain how LLMs build precise next token predictions. Another line of mechanistic interpretability works focus on inspecting factual knowledge encoded in the LLMs: they first locate such factual knowledge in pretrained LLMs Dai et al. (2022); Geva et al. (2023); Li et al. (2023b) and then edit them through updating a small set of parameters of LLMs Meng et al. (2022, 2023); Hase et al. (2023), which is so-called “locate-then-edit” model editing Ju and Zhang (2023). In this paper, we shed light on the mechanism of compositional reasoning on factual knowledge and borrow the idea from “locate-then-edit” model editing to locate the root and correct it to patch compositional reasoning failures of LLMs.
@@ D.1 Intervention
In Figure 4, the values is derived from the following procedures: (1) to make positive effect easier to observe, do clipping on the negative effects ($v<0\rightarrow v=0$) for each single datum; (2) do re-scaling to the effects for each single datum: for effects of different layers, min-max scale them to $[0,1]$; and (3) average the effects over the whole dataset.
@@ D.3 Important Module Locating
We implement our locating method on the basis of Causal Tracing Meng et al. (2022). Following  Meng et al. (2022)’s implementation, we also use a “window” intervention(a few layers before and after the intervened layer). In their original codebase, they set window size to be 10. In our experiments: we find that setting window size to be 2 is enough for us to effectively locate important modules. The locating results shown in Figure 5 adopted the window size of 6. In Figure 5, the values is derived from the following procedures: (1) do re-scaling to the indirect effects for each single datum: map the original value (before replacement) to 00 and map the positive indirect effects to $[0,1]$ and negative indirect effects to $[-1,0]$ in a linear manner; and (2) average the indirect effects over the whole dataset.
@@ D.4 Model Editing:CREME
The edit objective is the MHSA output matrix of $l$-th layer $W_{O}^{l}$. $W_{O}^{l}\in\mathbb{R}^{d\times d}$ operates as a key-value store for a set of vector keys $K=[k_{1}|k_{2}|...]$ and corresponding vector values $V=[v_{1}|v_{2}|...]$, by solving $(W_{O}^{l})^{T}K=V$. For a given compositional two-hop query and its corresponding second-hop query, we run the LLM twice: once with the compositional query and once with the second-hop query. We denote that: in the first pass with compositional query, the input of $W_{O}^{l}$ at the last token position is $k_{*}\in\mathbb{R}^{d\times 1}$; in the second pass with the corresponding second-hop query, the output of $W_{O}^{l}$ at the last token position is $v_{*}\in\mathbb{R}^{d\times 1}$. In practice, when calculating $k_{*}$ and $v_{*}$, we prepend tens of random tokens to the compositional query and the corresponding second-hop query to mimic context environments, and get multiple input vectors and output vectors. Then we average input vectors and output vectors of different context environment to get $k_{*}$ and $v_{*}$, respectively. The edited matrix is: $\hat{W_{O}^{l}}=W_{O}^{l}+(C^{-1}k_{*})^{T}\Lambda^{T}$ where $C=KK^{T}$ is a constant to estimate the uncentered covariance of $k$ (with a sample Wikipedia of text) and $\Lambda=(v_{*}-(W_{O}^{l})^{T}k_{*})/(C^{-1}k_{*})^{T}k_{*}$. In our experiments, we also apply a “window” editing (edit a few consecutive layers). In the main experiments, we adopt the window size of 6 in accordance with the locating experiments.
@@ E.6 Showcase of CREME
Method ROME (w. ground-truth) CREME (w.o. ground-truth)
Correction($\uparrow$) $\mathbf{98.0}\%$ $95.3\%$
Paraphrasing($\uparrow$) $62.5\%$ $\textbf{70.5}\%$
Generalization($\uparrow$) $+1.24$ $+\textbf{3.61}$
Specificity($\downarrow$) $+5.37$ $+\textbf{1.24}$
Input Types Paraphrasing $I_{p}$ Generalization $I_{g}$
$\Delta\log\text{PPL}$ $-29.9\%$ $-17.9\%$
Pred Accuracy
Pre-Patching $6.3\%$ $29.9\%$
Post-Patching $\mathbf{18.3}\%$ $\mathbf{38.3}\%$
Table 6: $\Delta\log\text{PPL}$ and Prediction Accuracy results for LLaMA-2-7B.